Please help and answer all

Calculate the molality of each solution.

Part A

0.495 mol solute; 2.17 kg solvent

molarity =_____m

Part B

0.559 mol solute; 1.44 kg solvent

molarity =_____m

Part C

0.139 mol solute; 538 g solvent

molarity =_____m
Answer #1

Molality = No. of moles of solute / weight of solvent in kg

Part A

0.495 mol solute; 2.17 kg solvent

Molality = (0.495 mol) / (2.17 kg) = 0.228 m

Part B

0.559 mol solute; 1.44 kg solvent

Molality = (0.559 mol) / (1.44 kg) = 0.39 m

Part C

0.139 mol solute; 538 g solvent

Molality  = (0.139 mol) / (0.538 kg) = 0.26 m